# thanks: Corresponding author

### v2

@@ 11            Sample random time steps $t$;
13            Compute $\mathcal{L}_{MSE}$ loss;
@@ IV.1.1 Datasets and Data preparation
In this study, we utilize three datasets and a test set, as described in Thuerey et al. (2020). These datasets are generated using RANS simulations to compute the velocity and pressure distributions of incompressible flows around airfoils. The simulations cover a wide range of Reynolds numbers ($Re=[0.5,5]$ million) and angles of attack (±$\pm 22.5^{\circ}$), providing a diverse set of aerodynamic conditions. A total of 1505 unique airfoil shapes were sourced from the UIUC database and combined with randomly sampled freestream conditions to create the training inputs. The RANS solutions were computed using the Spalart-Allmaras (SA) turbulence model, a widely used one-equation model for industrial applications. The computations were carried out with OpenFOAM, an open-source CFD software.
The first dataset is a regular dataset (Regular), which contains 220,000 images with a data size of (6, 128, 128). The six channels in each image represent: the initial velocity along the $x$-axis (channel 1), the initial velocity along the y-axis (channel 2), the airfoil mask (channel 3), pressure (channel 4), velocity along the $x$-axis (channel 5), and velocity along the $y$-axis (channel 6). The primary task of our approach is to train the model to correctly predict the pressure, $x$-axis velocity, and $y$-axis velocity based on the first three channels. The second dataset is an augmented dataset (Shear), which also consists of 220,000 images. It is generated by shearing the wing shape along the center axis by ±15 degrees to expand the shape space seen by the network, providing a richer variety of inputs. The final dataset is a small subset of the regular dataset (Subset), containing 6,000 images. This subset is introduced to better assess the model’s generalization ability when trained with a limited amount of data.
The test set comprises 90 images, which do not appear in the training data and remain unseen during all stages prior to the testing phase. The Reynolds number distribution in the test set spans a wide range, from [$2.04\times 10^{6}$ to $2.00\times 10^{7}$], encompassing 30 different airfoil types. All the datasets used in this study are publicly available TUM-I Repository 1 (2018); TUM-I Repository 2 (2018).
@@ IV.1.2 Configuration
We employ the DiT (Scalable Diffusion Models with Transformers Peebles and Xie (2023)) model as backbone for AeroDiT and adopt the same encoder and decoder models as in Ho, Jain, and Abbeel (2020). We use three configs: DiT-B, DiT-L and DiT-XL. Initially, we train two separate VAE-based encoder-decoder models for 100 epochs:
@@
After training, the conditional encoder is retained as the encoder, and the spatial decoder is retained as the decoder. The trained encoder and decoder are used to train the DiT model for 2,500,000 iterations. The experimental results are presented as the average of three random tests. During the testing phase, unless otherwise specified, the number of sampling steps is set to 100, and the model performance is evaluated using the L2 norm error between the predicted and ground truth.
Table 1: L2 Error of Airfoil Flow Simulations Across Multiple Datasets.
@@
Subset $0.1927^{\pm 0.23\%}$ $0.0262^{\pm 0.02\%}$ $0.0865^{\pm 0.03\%}$ $0.1989^{\pm 0.42\%}$ $0.0285^{\pm 0.14\%}$ $0.0909^{\pm 0.33\%}$ $0.1935^{\pm 0.13\%}$ $0.0271^{\pm 0.07\%}$ $0.0884^{\pm 0.05\%}$
Shear $0.2000^{\pm 0.39\%}$ $0.0470^{\pm 0.22\%}$ $0.1120^{\pm 0.30\%}$ $0.1964^{\pm 0.19\%}$ $0.0458^{\pm 0.13\%}$ $0.1116^{\pm 0.08\%}$ $0.1864^{\pm 0.76\%}$ $0.0415^{\pm 0.24\%}$ $0.0105^{\pm 0.37\%}$
Regular $0.1065^{\pm 0.58\%}$ $0.0251^{\pm 0.22\%}$ $0.0547^{\pm 0.22\%}$ $0.1041^{\pm 0.54\%}$ $0.0250^{\pm 0.23\%}$ $0.0537^{\pm 0.34\%}$ $0.1103^{\pm 0.26\%}$ $0.0278^{\pm 0.22\%}$ $0.0578^{\pm 0.31\%}$
@@ IV.2 Results and Discussion
Table 1 summarizes the prediction results of the model using different DiT backbones. In experiments conducted on the annotated dataset, the best predicted $L_{2}$ errors were consistently below 0.105, 0.025, and 0.055, respectively, across all test cases. The “$\pm$” values indicate the relative error percentage with respect to the mean $L_{2}$ error, providing a measure of variability across different test samples.
We provide more specific details on the model training pro- cess. Figure 4(a) illustrates the distribution of Reynolds numbers in the test set, which are uniformly distributed over a wide range from $2\times 10^{6}$ to $2\times 10^{7}$. Figure 4(b) shows the variation of the experimental loss with the number of training iterations for the backbone network, revealing that the loss stabilizes after approximately $5\times 10^{5}$ iterations. Finally, Figure 4(c) demonstrates that the model’s performance on the test domain becomes stable after more than $2\times 10^{5}$ training iterations.
@@ IV.3 Case Study
To evaluate the performance of AeroDiTs, we evaluated their performance by comparing predictions against CFD benchmarks across various airfoil geometries and Reynolds numbers, as shown in Table 3. Note that the selected airfoils and Reynolds numbers belong to the target domain, distinct from the source domain used for training. This ensures that the evaluation accurately reflects the model’s ability to generalize to unseen conditions. The chosen test cases are designed to assess the model’s generalization across a variety of airfoil shapes and a wide range of Reynolds numbers (2.04–20.00 million), which are representative of the complex turbulent flows encountered in real-world applications.
Table 3: Configuration of the airfoil and Reynolds number in the test case.
@@
Figure 5 shows the normalized pressure and velocity field obtained by AeroDiTs and the CFD of the Drela AG09 airfoil at $Re=4.04\times 10^{6}$. Qualitatively, the predictions of AeroDiTs are in good agreement with CFD results. For the pressure field, the maximum absolute error of the prediction result is around 0.09, while for the velocity field in both directions, the corresponding prediction error does not exceed 0.010 and 0.041. It can be found that the positions with the largest prediction errors all occur on the airfoil surface. Therefore, in order to quantitatively assess the prediction error, we give the flow field results along the airfoil surface, as shown in Figure 6. Overall, the longitudinal velocity along the airfoil surface is best predicted, with some differences in surface pressure and transverse velocity, but the overall trends of the curves are consistent.
The prediction errors on the airfoil surface are primarily concentrated in the pressure and transverse velocity fields, which are likely related to flow separation and boundary layer characteristics. In high Reynolds number flows, the turbulent structures and separation phenomena in the boundary layer significantly influence the flow field. AeroDiT may struggle with these complex flow features. From a physical perspective, these errors reflect the model’s limitations in capturing the fine details of the flow.
Moreover, AeroDiT achieves rapid predictions through its diffusion transformer architecture, significantly enhancing computational efficiency. However, this efficiency gain comes at the cost of some accuracy loss. While the model excels in predicting longitudinal velocity, errors persist in the pressure and transverse velocity predictions. This underscores the need to balance computational efficiency and prediction accuracy when designing deep learning models. For practical engineering applications, this trade-off should be adjusted according to specific needs. For instance, efficiency can be prioritized in the preliminary design phase, while higher accuracy is required in the optimization phase.
@@ V Discussion
Future directions could involve expanding the training dataset to include a broader range of high Reynolds number scenarios and incorporating physical laws into the learning process to enhance the model’s robustness and accuracy in more complex flow fields. Further research will also focus on developing domain generalization techniques Ma et al. (2023); Wang et al. (2022) based on varying Reynolds numbers. By incorporating precise Reynolds numbers as conditional inputs, we aim to further improve the predictive performance of the framework.
@@ VI Conclusion
This study introduces AeroDiT, a novel framework that applies Diffusion Transformers to make real-time, accurate flow field predictions. By leveraging DiTs’ ability to learn conditional distributions $p(\bm{x}|\bm{\Psi})$, AeroDiT achieves reliable predictions of pressure and velocity fields, with average L2 errors of 0.1 for pressure ($p$) and 0.03 and 0.05 for velocity components ($u_{x}$, $u_{y}$), respectively. The results demonstrate the model’s ability to generalize across a broad range of Reynolds numbers ($2.04\times 10^{6}$ to $2.00\times 10^{7}$) and diverse airfoil geometries. In test cases involving the Drela AG09 and AH 63-K-127/24 airfoils, AeroDiT shows excellent agreement with CFD benchmarks, with prediction errors for pressure and velocity fields consistently below 0.02, even under challenging conditions.
Notably, AeroDiT exhibits strong performance in near-real-time scenarios, with the sampling steps converging to stable predictions within seconds. The transformer-based architecture significantly reduces computational costs, making high Reynolds number aerodynamic simulations more feasible for real-time applications. However, in cases involving extreme Reynolds numbers ($>1.2\times 10^{7}$), the model’s performance shows some limitations, likely due to the scarcity of corresponding training data and the drastic changes in flow field characteristics at these scales.
These findings highlight AeroDiT’s potential to bridge the gap between computational efficiency and predictive accuracy in CFD simulations. Looking ahead, the continued development and refinement of AeroDiT hold the promise of even greater advancements. By enriching the training dataset with higher Reynolds number scenarios and integrating physical constraints, we can further enhance the model’s performance in complex turbulence regimes. This not only paves the way for more accurate simulations but also opens new avenues for exploring the intricate dynamics of airflow around airfoils.
As we stand on the cusp of a new era in computational fluid dynamics, AeroDiT stands as a testament to the transformative power of machine learning. Its potential applications in various fields, from aerospace engineering to renewable energy, are vast and far-reaching. By bridging the gap between computational efficiency and predictive accuracy, AeroDiT heralds a future where real-time aerodynamic simulations are not just a possibility but a reality.
Appendix A ref figures
Figure 7: Distribution of pressure and velocity fields around the airfoil from low to high Reynolds numbers for the same airfoil : $Re_{1}=9.11\times 10^{6}$, $Re_{2}=1.06\times 10^{7}$, $Re_{3}=2.00\times 10^{7}$
Figure 8: Distribution of pressure and velocity fields around different airfoils at the same high Reynolds number : $Re=1.21\times 10^{7}$.

### v3

@@ 11     Sample random time steps $t$;
13     Compute $\mathcal{L}_{\text{total}}$;
@@ IV.1.1 Datasets and Data preparation
In this study, we utilize three datasets and a test set, as described in Thuerey et al. (2020). These datasets are generated using RANS simulations to compute the velocity and pressure distributions of incompressible flows around airfoils. The simulations cover a wide range of Reynolds numbers ($Re=[0.5,5]$ million) and angles of attack ($\pm 22.5^{\circ}$), providing a diverse set of aerodynamic conditions. A total of 1505 unique airfoil shapes were sourced from the UIUC database and combined with randomly sampled freestream conditions to create the training inputs. The RANS solutions were computed using the Spalart-Allmaras (SA) turbulence model, a widely used one-equation model for industrial applications. The computations were carried out with OpenFOAM, an open-source CFD software.
The first dataset is a regular dataset (Regular), which contains 220,000 images with a data size of (6, 128, 128). The six channels in each image represent: the initial velocity along the $x$-axis (channel 1), the initial velocity along the y-axis (channel 2), the airfoil mask (channel 3), pressure (channel 4), velocity along the $x$-axis (channel 5), and velocity along the $y$-axis (channel 6). The primary task of our approach is to train the model to correctly predict the pressure, $x$-axis velocity, and $y$-axis velocity based on the first three channels. The second dataset is an augmented dataset (Shear), which also consists of 220,000 images. It is generated by shearing the wing shape along the center axis by $\pm$15 degrees to expand the shape space seen by the network, providing a richer variety of inputs. The final dataset is a small subset of the regular dataset (Subset), containing 6,000 images. This subset is introduced to better assess the model’s generalization ability when trained with a limited amount of data.
The test set comprises 90 images, which do not appear in the training data and remain unseen during all stages prior to the testing phase. The Reynolds number distribution in the test set spans a wide range, from $2.04\times 10^{6}$ to $2.00\times 10^{7}$], encompassing 30 different airfoil types. All the datasets used in this study are publicly available TUM-I Repository 1 (2018); TUM-I Repository 2 (2018).
@@ IV.1.2 Configuration
We employ the DiT (Scalable Diffusion Models with Transformers Peebles and Xie (2023)) model as the backbone for AeroDiT and adopt the same encoder and decoder models as in Ho, Jain, and Abbeel (2020). We use three configurations: DiT-B/2, DiT-L/2, and DiT-XL/2. The suffix "/2" indicates that the model operates on a latent feature grid that is downsampled by a factor of 2 in spatial resolution, achieved through a patch size of $2\times 2$ during tokenization.
Initially, we train two separate VAE-based encoder-decoder models for 100 epochs:
@@
After training, the conditional encoder is retained as the encoder, and the spatial decoder is retained as the decoder. The trained encoder and decoder are then used to train the DiT model for 2,500,000 iterations. The model is trained using a global batch size of 4 for 150 epochs with the AdamW optimizer (learning rate of $10^{-4}$, no weight decay). Unless otherwise specified, the sampling process during inference uses 100 steps.
Following previous works, we evaluate model performance using the mean relative $L_{2}$ error, defined as:
$\text{Relative }L_{2}\text{ Error}=\frac{\|\hat{\bm{x}}-\bm{x}\|_{2}}{\|\bm{x}\|_{2}},$ (13)
where $\hat{\bm{x}}$ is the predicted field and $\bm{x}$ is the ground truth.
We also report the maximum absolute error in selected case studies, defined as:
$\text{Max Abs Error}=\max_{i}|\hat{x}_{i}-x_{i}|.$ (14)
We adopt the standard DDPM sampling procedure during inference, implemented via p_sample_loop(). Although our codebase also includes DDIM-based alternatives, we found that DDPM offers slightly better accuracy and more stable predictions in our fluid dynamics setting. Therefore, DDPM is used consistently throughout all experiments reported in this paper.
For completeness, we list the architecture details of the DiT variants used in this work:
  • DiT-B/2: 12 transformer blocks, 768 hidden size, 12 attention heads, patch size = 2 (i.e., $64\times 64$ token map).
  • DiT-L/2: 24 transformer blocks, 1024 hidden size, 16 attention heads, patch size = 2.
  • DiT-XL/2: 28 transformer blocks, 1280 hidden size, 16 attention heads, patch size = 2.
The patch size of 2 means the input latent tensor of size $128\times 128$ is divided into $64\times 64$ non-overlapping patches, each of size $2\times 2$. Each patch is then embedded and processed as a token within the DiT transformer backbone.
Table 1: Mean Relative $L_{2}$ Error Across Multiple Datasets.
@@
Subset $1927^{\pm 0.23\%} 0.0$$262^{\pm 0.02\%} 0.0$$0865^{\pm 0.03\%}$ $1989^{\pm 0.42\%} 0.0$$285^{\pm 0.14\%} 0.0$$0909^{\pm 0.33\%}$ $1935^{\pm 0.13\%} 0.0$$271^{\pm 0.07\%} 0.0$$0884^{\pm 0.05\%}$
Shear $2000^{\pm 0.39\%} 0.0$$0470^{\pm 0.22\%}$ $0.1120^{\pm 0.30\%}$ $1964^{\pm 0.19\%} 0.0$$0458^{\pm 0.13\%}$ $0.1116^{\pm 0.08\%}$ $1864^{\pm 0.76\%} 0.0$$415^{\pm 0.24\%} 0.0$$0105^{\pm 0.37\%}$
Regular $1065^{\pm 0.58\%} 0.0$$251^{\pm 0.22\%} 0.0$$0547^{\pm 0.22\%}$ $1041^{\pm 0.54\%} 0.0$$250^{\pm 0.23\%} 0.0$$0537^{\pm 0.34\%}$ $1103^{\pm 0.26\%} 0.0$$278^{\pm 0.22\%} 0.0$$0578^{\pm 0.31\%}$
Figure 4: Visualization of input data (Airfoil geometry, initial horizontal velocity, initial parallel velocity), prediction (Pressure Field, velocity X, velocity Y) and the corresponding ground truth.
@@ IV.2 Results and Discussion
Table 1 summarizes the prediction results of the model using different DiT backbones. In experiments conducted on the annotated dataset, the best predicted $L_{2}$ errors were consistently below $0.105$, $0.025$, and $0.055$, respectively, across all test cases. Beyond $p,u_{x},u_{y}$, we also report turbulence-relevant metrics (spanwise vorticity $\omega_{z}$, kinetic energy, enstrophy), computed via $3{\times}3$ centered differences over the fluid region; on the 90-case test set the mean relative $L_{2}$ errors are $0.0777$ ($\omega$), $0.0399$ (KE), and $0.0764$ (enstrophy).
Figure 4 presents the performance of the DiT-L/2 model trained on the Regular dataset for three different airfoil configurations, corresponding to Reynolds numbers of $2.04\times 10^{6}$, $1.06\times 10^{7}$, and $2.00\times 10^{7}$, representing the minimum, median, and maximum Reynolds numbers in the test set. Additionally, the differences between the predicted and ground-truth values are visualized. The discrepancies between predictions and ground truth are minimal and nearly indistinguishable to the naked eye. Furthermore, the model’s performance remains consistently reliable across both high and low Reynolds numbers, with negligible variation.
We provide more specific details on the model training process. Figure 5(a) illustrates the distribution of Reynolds numbers in the test set, which are uniformly distributed over a wide range from $2\times 10^{6}$ to $2\times 10^{7}$. Figure 5(b) shows the variation of the experimental loss with the number of training iterations for the backbone network, revealing that the loss stabilizes after approximately $5\times 10^{5}$ iterations. Finally, Figure 5(c) demonstrates that the model’s performance on the test domain becomes stable after more than $2\times 10^{5}$ training iterations.
@@ IV.3 Case Study
To evaluate the performance of AeroDiT, we compare predictions against CFD benchmarks across various airfoil geometries and Reynolds numbers, as shown in Table 3. Note that the selected airfoils and Reynolds numbers belong to the target domain, distinct from the source domain used for training. This ensures that the evaluation accurately reflects the model’s ability to generalize to unseen conditions. The chosen test cases assess generalization across diverse airfoil shapes and a wide range of Reynolds numbers (2.04–20.00 million), representative of the complex turbulent flows encountered in practical applications.
Table 3: Configuration of the airfoil and Reynolds number in the test cases.
@@
Figure 6 shows the normalized pressure and velocity fields predicted by AeroDiT and the CFD solution for the Drela AG09 airfoil at $Re=4.04\times 10^{6}$. Qualitatively, AeroDiT is in good agreement with CFD. For the pressure field, the maximum absolute error is around 0.09, while for the velocity field in both directions, the corresponding errors do not exceed 0.010 and 0.041. We observe that the largest errors concentrate along the airfoil surface. To quantify these, Figure 7 compares pressure and velocity along the surface: the streamwise velocity is captured best, whereas differences appear in surface pressure and transverse velocity. Physically, these residuals likely stem from boundary-layer behavior and potential local separation at high Reynolds numbers, where fine-scale turbulent structures dominate near-wall dynamics.
Figure 8: Normalized flow field of AH 63-K-127/24 airfoil at $Re=9.11\times 10^{6}$.
Figure 9: Comparison of surface pressure and velocities of AH 63-K-127/24 airfoil at $Re=9.11\times 10^{6}$.
In test case 2, the airfoil is changed to AH 63-K-127/24 and the Reynolds number is increased to $9.11\times 10^{6}$. The normalized flow fields are shown in Figure 8. Somewhat counterintuitively, AeroDiT performs even better than in test case 1: absolute errors are mostly below 0.075 (pressure), 0.020 ($u_{x}$), and 0.045 ($u_{y}$). The surface comparisons in Figure 9 confirm the improved curve match. A plausible explanation is that the training data distribution contains more samples in the vicinity of this Reynolds-number regime (order $10^{7}$), which improves conditional coverage and reduces extrapolation. From a physics perspective, the boundary layer remains attached over a larger portion of the chord in this case, mitigating near-wall discrepancies.
Figure 10: Normalized flow field of EPPLER 59 airfoil at $Re=1.21\times 10^{7}$.
Figure 11: Comparison of surface pressure and velocities of EPPLER 59 airfoil at $Re=1.21\times 10^{7}$.
Finally, the airfoil is EPPLER 59 at $Re=1.21\times 10^{7}$. The flow-field and surface-quantity comparisons are shown in Figures 10 and 11. Here the model performance degrades, especially in surface pressure and transverse velocity. Two factors likely contribute: (i) fewer high-$Re$ cases at this order in the training set, and (ii) pronounced changes in flow physics when moving from $O(10^{6})$ to $O(10^{7})$, including stronger adverse-pressure-gradient effects and earlier separation onset. These observations suggest that adding more high-$Re$ training samples and/or strengthening physics constraints (e.g., RANS-residual losses) could further improve robustness in this regime.
@@ V Discussion
Beyond predictive accuracy, one notable advantage of AeroDiT lies in its inference efficiency. While diffusion transformers are indeed computationally intensive during training, our model only requires a single training run to generalize across diverse airfoil shapes and flow conditions. Once trained, AeroDiT generates physically consistent flow fields within seconds on a single GPU, which is orders of magnitude faster than traditional CFD solvers such as RANS. This enables rapid design evaluation and real-time inference in practical applications, demonstrating a substantial computational gain in engineering workflows.
@@
Future directions could involve expanding the training dataset to include a broader range of high Reynolds number scenarios to enhance the model’ robustness and accuracy in more complex flow fields. Further research will also focus on developing domain generalization techniques Ma et al. (2023); Wang et al. (2022) based on varying Reynolds numbers. By incorporating precise Reynolds numbers as conditional inputs, we aim to further improve the predictive performance of the framework.
@@ VI Conclusion
This study introduces AeroDiT, a novel framework that applies Diffusion Transformers to make real-time, accurate flow field predictions. By leveraging DiTs’ ability to learn conditional distributions $p(\bm{x}|\bm{\Psi})$, AeroDiT achieves reliable predictions of pressure and velocity fields, with average L2 errors of 0.1 for pressure ($p$) and 0.03 and 0.05 for velocity components ($u_{x}$, $u_{y}$), respectively. The results demonstrate the model’ ability to generalize across a broad range of Reynolds numbers ($2.04\times 10^{6}$ to $2.00\times 10^{7}$) and diverse airfoil geometries. In test cases involving the Drela AG09 and AH 63-K-127/24 airfoils, AeroDiT shows excellent agreement with CFD benchmarks, with prediction errors for pressure and velocity fields consistently below 0.02, even under challenging conditions.
Notably, AeroDiT exhibits strong performance in near-real-time scenarios, with the sampling steps converging to stable predictions within seconds. The transformer-based architecture significantly reduces computational costs, making high Reynolds number aerodynamic simulations more feasible for real-time applications. However, in cases involving extreme Reynolds numbers ($1.2\times 10^{7}$), the model’ performance shows some limitations, likely due to the scarcity of corresponding training data and the drastic changes in flow field characteristics at these scales.
In addition, we incorporate explicit physics-informed loss terms based on RANS residuals, including mass and momentum conservation constraints. This integration improves the physical plausibility of predictions and enhances training stability, further supporting AeroDiT as a hybrid physics-ML modeling framework.
These findings highlight AeroDiT’ potential to bridge the gap between computational efficiency and predictive accuracy in CFD simulations. Future work should focus on enriching the training dataset with higher Reynolds number scenarios and extending the model to unsteady and three-dimensional flow conditions, as well as exploring more complex physics constraints.